OpenStudy (anonymous):

please help ...an arrow is shot vertically upward with aspeed of 288ft per sec, and 3s later, another is shot up at speed of 240ft per sec.will they meet? if so where??

OpenStudy (anonymous):

when they meet, their displacement is the same. \(\vec S=\vec u t + \frac{1}{2}\vec a t^2 \) \(\vec v^2 = \vec u^2 +2\vec a \cdot \vec S\) sub it all in and solve note: one of them will have moved for a time t and the other, t+3 do you need further help?